# LexEval: A Comprehensive Chinese Legal Benchmark for Evaluating Large Language Models

### v2

@@ 2 Related Work
However, despite great success, they face significant challenges in the legal domain [31, 7, 14, 30]. In the legal domain, accuracy, reliability, and fairness are crucial, but LLMs often perform poorly in these aspects due to issues like hallucination [31] and inherent biases [51, 9]. Hallucination refers to models generating information that is not based on facts, which can lead to misleading or entirely incorrect conclusions in legal documents and consultations. Additionally, due to biases in the training data, the model may inadvertently replicate and amplify these biases, affecting its fairness and accuracy in applications such as legal judgment prediction, case analysis, and contract review.
@@ 4.1 Setup
We evaluate the LLMs in both zero-shot and few-shot settings. In the zero-shot setting, the inputs to LLMs are only instructions and queries. In the few-shot setting, we design three different examples for each task. These examples can be found on the GitHub website. When evaluating LLMs, we set the temperature to 0 to minimize the variance introduced by random sampling. For chat LLMs, we reserve the format of their dialog prompts. When the input length exceeds the maximum context length of LLMs, we truncate the input sequence from the middle since the front and end of the input may contain crucial information. The input prompts used during our evaluation can be found in the Appendix C.4. We standardize our evaluation metrics by using Accuracy to evaluate all multiple-choice questions and Rough-L to evaluate tasks at Generation level. The evaluation metrics for each task can be found in Table  1. We also discuss the limitations of the evaluation metrics in Appendix  B.1.
@@ 4.2 Evaluated Models
ChatGLM2 19.8 21.1 28.3 15.0 24.1 14.0 36.8 27.2 52.2 29.8 10
Chinese-Alpaca-2-7B 27.8 24.7 28.6 15.7 31.2 14.6 21.5 28.4 40.4 29.4 11
ChatLaw-33B 10.0 17.1 21.4 7.0 14.2 13.2 15.3 19.1 34.2 28.7 12
@@ C.4 Task Instruction and example
In this section, we present the task Instruction for each task. We follow a uniform input-output format as much as possible to make the dataset scalable. Table 7 through Table 29 provide illustrative examples for each task category. Specifically, Tables 7 to 9 exemplify tasks at Memorization level, while Tables 10 to 14 showcase Understanding tasks. Logic Inference tasks are exemplified in Tables 15 to 20, and Discrimination tasks are illustrated in Tables 21 and 22. Generation tasks are represented by Tables 23 to 26, and Ethic tasks are demonstrated in Tables 27 to 29.
@@ Appendix E More Evaluation Result
Due to the length limitations of the paper, a series of specific results are not fully presented. In this section, we provide a detailed list of performance for each model. Specifically, Tables  31 and  32 show the performance in the zero-shot setting. Tables  33 and  34 demonstrate the performance in the few-shot setting. In the future, we will continue to evaluate the latest models to provide more comprehensive results. All evaluation experiments were conducted on an Ubuntu server equipped with a 128-core Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 8 NVIDIA A100 SXM 80GB GPUs. Additionally, the CUDA version was 11.7, the Python version was 3.9.0, the PyTorch version was 2.0.0, and the transformers version was 4.28.1.

### v3

@@ 2 Related Work
However, despite great success, they face significant challenges in the legal domain [32, 7, 14, 31]. In the legal domain, accuracy, reliability, and fairness are crucial, but LLMs often perform poorly in these aspects due to issues like hallucination [32] and inherent biases [52, 9, 29]. Hallucination refers to models generating information that is not based on facts, which can lead to misleading or entirely incorrect conclusions in legal documents and consultations. Additionally, due to biases in the training data, the model may inadvertently replicate and amplify these biases, affecting its fairness and accuracy in applications such as legal judgment prediction, case analysis, and contract review.
@@ 4.1 Setup
We evaluate the LLMs in both zero-shot and few-shot settings. In the zero-shot setting, the inputs to LLMs are only instructions and queries. In the few-shot setting, we design three different examples for each task. These examples can be found on the GitHub website. When evaluating LLMs, we set the temperature to 0 to minimize the variance introduced by random sampling. For chat LLMs, we reserve the format of their dialog prompts. When the input length exceeds the maximum context length of LLMs, we truncate the input sequence from the middle since the front and end of the input may contain crucial information. The input prompts used during our evaluation can be found in the Appendix C.4. We standardize our evaluation metrics by using Accuracy to evaluate all multiple-choice questions and Rough-L to evaluate tasks at Generation level.
The evaluation metrics for each task can be found in Table  1. We also discuss the limitations of the evaluation metrics in Appendix  B.1.
@@ 4.2 Evaluated Models
ChatLaw-33B 10.0 17.1 23.8 9.9 15.2 13.3 15.3 19.1 34.2 30.0 10
ChatGLM2 20.2 21.1 28.4 15.5 24.1 14.0 36.8 27.2 52.2 29.9 11
Chinese-Alpaca-2-7B 27.8 24.7 28.6 15.7 31.2 14.6 21.5 28.4 40.4 29.4 12
@@ 5 Conclusion & Future Work
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix B.1.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix B.2.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Section 3.4.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Appendix A.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] We don’t have a training process.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Because running a massive amount of LLMs is expensive.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 3.2.
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix C.1.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See Appendix A.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix C.1.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix 3.4.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.2, Appendix C.4 and Appendix F.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 3.2.
@@ C.4 Task Instruction and example
In this section, we present the task Instruction for each task. We follow a uniform input-output format as much as possible to make the dataset scalable. Table 8 through Table 30 provide illustrative examples for each task category. Specifically, Tables 8 to 10 exemplify tasks at Memorization level, while Tables 11 to 15 showcase Understanding tasks. Logic Inference tasks are exemplified in Tables 16 to 21, and Discrimination tasks are illustrated in Tables 22 and 23. Generation tasks are represented by Tables 24 to 27, and Ethic tasks are demonstrated in Tables 28 to 30.
Table 7: Overview of LexEval in Comparison with Existing General and Legal Domain Benchmarks.
Dataset Language Domain Data Source Taxonomy Task Num Data Size
C-Eval_Legal Chinese General existing datasets - 2 493
CMMLU_Legal Chinese General existing datasets - 1 216
LEAGALBENCH English Legal
existing datasets
expert annotation
issue-spotting
rule-recall
rule-application
rule-conclusion
162 -
LawBench Chinese Legal existing datasets
Memorization
Understanding
Applying
20 10000
LAiW Chinese Legal existing datasets
basic information retrieval
legal foundation inference
complex legal application
14 11605
LexEval Chinese Legal
existing datasets,
exam,
expert annotation
Memotization
Understanding
Logic Inference
Discrimination
Generation
Ethic
23 14150
C.5 Comparison with Existing Benchmarks
As detailed in Table 7, we have compared existing legal benchmarks—both general and Chinese law-specific—across several key criteria: language, domain, data source, taxonomy, number of tasks, and data size. C_eval and CMMLU are general-domain datasets that include a limited subset of legal evaluation data. We have only included this portion in Table 7. LEGALBENCH is an English legal benchmark comprising 162 tasks contributed by 40 contributors. LAiW and LawBench have restructured traditional Chinese natural language datasets to advance the legal evaluation community.
@@ Appendix E More Evaluation Result
Due to the length limitations of the paper, a series of specific results are not fully presented. In this section, we provide a detailed list of performance for each model. Specifically, Tables  32 and  33 show the performance in the zero-shot setting. Tables  34 and  35 demonstrate the performance in the few-shot setting. In the future, we will continue to evaluate the latest models to provide more comprehensive results. All evaluation experiments were conducted on an Ubuntu server equipped with a 128-core Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 8 NVIDIA A100 SXM 80GB GPUs. Additionally, the CUDA version was 11.7, the Python version was 3.9.0, the PyTorch version was 2.0.0, and the transformers version was 4.28.1.